# Critical Sections Are Not Per-Thread: A Trace Semantics for Lock-Based Concurrency

### 20 Mar 2026 v2

@@ 4. Related Work
Our results apply independently of how traces are obtained (e.g., via instrumentation, symbolic execution, or CFG reachability). Static analysis works typically consider the reachability graph. There are numerous works that employ the lock set method for static deadlock detection, e.g., see [KPSW16, Sha08, NPSG09, NPSG09, CYSZ22, KPSW16].
@@ 5. Conclusion
Pretty much all prior works assume per-thread lock sets and critical sections restricted to a single thread. This assumption does not hold for general C/Pthread executions. Critical sections that cover multiple threads exist, can be observed in simple programs and invalidate the standard foundation. Similar observations apply to Java as well. Correcting this semantic gap is important because lock sets form the foundation of numerous analyses for deadlock and data race detection. We consider a minimal trace model for C/Pthread programs and give corrected definitions based on a trace-based characterization of critical sections and lock sets to deal with programs where locks protect events across thread boundaries. Extending the semantics with memory operations and additional synchronization primitives such as condition variables introduces orthogonal challenges and is left for future work.

### 07 Apr 2026 v4

@@ 4. Related Work
Our results apply independently of how traces are obtained (e.g., via instrumentation, symbolic execution, or CFG reachability). Static analysis works typically consider the reachability graph. There are numerous works that employ the lock set method for static deadlock detection, e.g., see [KPSW16, Sha08, NPSG09, CYSZ22, KPSW16].
@@ 5. Conclusion
Pretty much all prior works assume per-thread lock sets and critical sections restricted to a single thread. This assumption does not hold for general C/Pthread executions. Critical sections that cover multiple threads exist, can be observed in simple programs and invalidate the standard foundation. Similar phenomena can occur in other thread-and-monitor models such as Java. Correcting this semantic gap is important because lock sets form the foundation of numerous analyses for deadlock and data race detection. We consider a minimal trace model for C/Pthread programs and give corrected definitions based on a trace-based characterization of critical sections and lock sets to deal with programs where locks protect events across thread boundaries. Extending the semantics with memory operations and additional synchronization primitives such as condition variables introduces orthogonal challenges and is left for future work.